Find the distance between the points R(-4,1) and S(2,7)
Answer by TimothyLamb(4379)

use the pythagorean theorem on the square of the differences in x and y
---
d = sqrt( (2+4)^2 + (7-1)^2 )
d = sqrt( (6)^2 + (6)^2 )
d = sqrt( 36 + 36 )
d = sqrt( 72 )
---
d ~= 8.485
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php